# The role of three-dimensional effects on ion injection and acceleration in perpendicular shocks

### v1

@@ 4 Conclusions
By varying the shock Mach number, we show that higher $M_{A}$ shocks produce harder spectra and more efficient particle acceleration. Notably, our high-resolution simulations suggest that the threshold Mach number required to achieve a canonical $p^{-4}$ spectrum is likely underestimated in low-resolution studies (see Paper I). This reinforces the importance of both high spatial resolution and large simulation domains for properly capturing the nonlinear dynamics of perpendicular shocks. For oblique shocks, which are not covered in this work, the initial acceleration mechanism would still involve SDA. However, in this case, particles could escape upstream by streaming along the magnetic field lines, which is not possible in purely perpendicular shocks. The resulting escaping current could amplify the magnetic field, triggering a second phase of evolution where particles transition from SDA to DSA. This scenario will be explored in a future paper.
L.O. thanks Siddhartha Gupta and Robert Ewart for insightful discussions. Simulations were performed on computational resources provided by the University of Chicago Research Computing Center. L.O. acknowledges the support of the Multimessenger Plasma Physics Center (MPPC), NSF grants PHY2206607. D.C. was partially supported by NASA through grants 80NSSC20K1273 and 80NSSC18K1218 and NSF through grants AST-1909778, PHY-2010240, and AST- 2009326. L.S. was supported by NSF grant PHY2409223. This research was facilitated by Multimessenger Plasma Physics Center (MPPC) NSF grants PHY2206607 and PHY2206609 to A.S. and L.S. The work was supported by a grant from the Simons Foundation (MP-SCMPS-0000147, to L.S. and A.S.). L.S. also acknowledges support from DoE Early Career Award DE-SC0023015.

### v2

@@ IV Conclusions
By varying the shock Mach number, we show that higher $M_{A}$ shocks produce harder spectra and more efficient particle acceleration. Notably, our high-resolution simulations suggest that the threshold Mach number required to achieve a canonical $p^{-4}$ spectrum is likely underestimated in low-resolution studies (see Paper I). This reinforces the importance of both high spatial resolution and large simulation domains for properly capturing the nonlinear dynamics of perpendicular shocks. We chose an angle of 90° in order to study the extreme, purely perpendicular limit, where the background magnetic field has no component along the shock normal. This configuration allows us to obtain a clean and controlled result, isolating the effects of perpendicular shocks without interference from parallel-field dynamics, and providing a clear benchmark for understanding particle acceleration and field amplification in this limit. For oblique shocks, which are not covered in this work, the initial acceleration mechanism would still involve SDA. However, in this case, particles could escape upstream by streaming along the magnetic field lines, which is not possible in purely perpendicular shocks. The resulting escaping current could amplify the magnetic field, triggering a second phase of evolution where particles transition from SDA to DSA. This scenario will be explored in a future paper.
L.O. thanks Siddhartha Gupta and Robert Ewart for insightful discussions. Simulations were performed on computational resources provided by the University of Chicago Research Computing Center. L.O. acknowledges the support of the Multimessenger Plasma Physics Center (MPPC), NSF grant PHY2206607. D.C. was partially supported by NASA through grants 80NSSC20K1273 and 80NSSC18K1218 and NSF through grants AST-1909778, PHY-2010240, and AST- 2009326. L.S. was supported by NSF grant PHY2409223. This research was facilitated by Multimessenger Plasma Physics Center (MPPC) NSF grants PHY2206607 and PHY2206609 to A.S. and L.S. The work was supported by a grant from the Simons Foundation (MP-SCMPS-0000147, to L.S. and A.S.). L.S. also acknowledges support from DoE Early Career Award DE-SC0023015.